# Learning in Wilson-Cowan model for metapopulation

### v1

@@ 3 Enforcing Linear Stability
the maximum value between the two real part of (7), i.e., $\max\{\text{Re}(s^{(+)}),\text{Re}(s^{(-)})\}$, must be taken. If such a value is negative, then the associated fixed point is asymptotically stable (Strogatz, 2018).
To be explicit, the elements of $[\mathbf{J}]_{ii}(\vec{\mathbf{z}}^{*})$ have the following expressions: {strip}
@@ 4 Training
Obtained the stability criteria for our attractors, we can construct a statistical learning model for our classification task. In such a statistical learning model only the components of the matrix $\Phi$, which are not the embedded eigenvectors, the non-zero eigenvalues in $\Lambda$, and the $\gamma$ parameter can be learned. In other words, the matrix $\Phi$ should be regarded as a matrix where the first $K$ columns, identifying the embedded attractors of the system’s dynamics, are not to be learned, just as the first $K$ eigenvalues of the matrix $\Lambda$, which are fixed to be zero by construction. This approach ensures that the planted attractors guide the dynamics, letting the system to converge on selected attractors once the training has been completed. As an initial condition for the non-embedded eigenvectors, we opt for orthogonal random eigenvectors (Bansal et al., 2018). As an initial condition for $\gamma$ we opt for a value equal to $0.25$ (as described in Section 2), while for the trainable eigenvalues we opt to initialize them with a normal distribution with mean equal to $-\sqrt{\mathcal{N}}$ and variance equal to $1$. The condition on the trainable eigenvalues satisfies completely the initial stability condition (see Fig. 3). Without loss of generality, we fix the non trainable parameters of the model, as described in the caption of Fig. 3 and used in Zankoc et al. (2017).
@@
Figure 4: Building the model. The figure illustrates the composition of the dataset and the method for introducing stable attractors into the dynamics. Panel a) displays the labels, images, and targets used in our model. The first row shows the labels of the MNIST handwritten image examples found in the second row. The third row presents the ten target images we created to serve as attractors for the dynamics. Panel b) shows the creation of an eigenvector of matrix $\mathbf{A}$. The image is composed solely of the excitatory stationary parts of the two fixed points in the Wilson-Cowan model. Each black pixel corresponds to a $\overline{x}^{(1)}$, while each white pixel corresponds to a $\overline{x}^{(2)}$. Panel c) illustrates the positions of the fixed eigenvectors and eigenvalues (in red) and the trainable ones (in green). Recall that the matrix $\mathbf{A}$ is diagonalizable by definition.
@@
The input datum $\mathbf{z}^{(j)}$ represents, for our dynamical system defined in (1), an initial condition of the dynamics, i.e., $\mathbf{z}^{(j)}(t=0)$. From such initial condition, we evolve the dynamical system for a sufficiently large time $T$. The evolution is performed using Euler’s algorithm (Griffiths and Higham, 2010), discretizing time into steps of $\Delta t=0.1$, unless explicitly specified otherwise. At each iteration, denoted with the index $n$, the system’s state $\mathbf{z}^{(j)}_{n}$ is updated according to the equation: $\mathbf{z}^{(j)}_{n+1}=\mathbf{z}^{(j)}_{n}+G(\mathbf{z}^{(j)}_{n})\Delta t$, where $G(\mathbf{z}^{(j)})$ is the system’s flux, defined in (1). The total number of iterations, $n_{max}$, is determined a priori based on the desired simulation time $T$ and the chosen time step $\Delta t$ through the relation $n_{max}=\lfloor\frac{T}{\Delta t}\rfloor.$ This approach allowed us to efficiently simulate the system’s long-term behavior by iteratively applying Euler’s update rule. Once the system reaches the final iteration, we use the value of the solution to optimize the loss function.
To minimize the loss function in (9), we specifically employ Accelerated Stochastic Gradient Descent using the Adam optimizer (Kingma and Ba, 2015), with a learning rate of $0.1$, unless explicitly specified otherwise. The learning rate, the number of epochs needed to achieve optimal minimization of the loss function, the mini-batch size used in the Adam algorithm, and the time $T$ required to reach the stationary state have been optimized empirically (see A).
@@
Figure 5: Dynamics of the model. The figure illustrates a schematic representation of our Neural Mass Network Model. Panel a) displays an image from the MNIST dataset along with its associated target. The image features a black pixel with a red border, which we track throughout its dynamics. Panel b) shows this dynamic process. The equations are highlighted in red to emphasize that we are focusing on that single pixel. This pixel is depicted as a red sphere moving within a double well potential. Our model, we recall, is configured to exhibit a bistable profile. Initially, the vector $\vec{\zeta}_{i}$ is set to represent the black pixel for both subpopulations of neurons. As the dynamics evolve according to our model’s equations, the pixel explores the bistable profile. It does not stop at the closest minimum; instead, it reaches the correct minimum designated for classification, thanks to the learned coupling in the matrix $\mathbf{A}$. Panel c) illustrates the overall dynamics of the entire image. Starting from an initial condition—the image we want to classify—the system uses the equations in (1) to achieve the final state, which is our stable attractor. The precision cutoff for this analysis has been set to the seventh decimal digit.
@@
$\delta(\arg\max\big{(}\vec{q}^{(j)}\big{)})$ is, therefore, a binary random variable that allows us to estimate the probability of success of our algorithm using a simple Horvitz-Thompson estimator (Horvitz and Thompson, 1952), i.e., equation (11).
@@ 5 Experimental Validations
    Model Fine-Tuning with Pre-Trained Weights: The second subsection explores the test set performance of a Wilson-Cowan model for metapopulation in combination with convolutional neural networks (Gu et al., 2018) or a Pre-Trained Transformer (Vaswani et al., 2017a). In the first case we will analyse the performance of the combination across MNIST, Fashion-MNIST, CIFAR10, and TF-FLOWERS. While in the second case, we perform a classification for the sentiment analysis, i.e. IMDB dataset (Pal et al., 2020). For details on these datasets, see A.
@@ 5.1 Training from Scratch
A critical distinction between the top panel of Fig. 6 and Fig. 3 lies in the value of the parameter $\gamma$, which plays a key role in governing the time scales within the standard Wilson-Cowan model. While traditionally it is set to $0.25$, our work explores the impact of optimizing $\gamma$ during the training process. In Fig. 3, $\gamma$ begins with its typical value of $0.25$. As stated before, a population in each vertex is composed by $80\%$ of excitatory and $20\%$ inhibitory neurons. However, during training, it undergoes optimization and converges to an optimal value of $0.839\pm 0.031$ for the MNIST dataset ( for Fashion MNIST results see Tab. 1). In this case, we have that a population in a vertex is composed by $54\%$ of excitatory and $46\%$ inhibitory neurons. These values suggest that, for our particular Neural Mass Network Model, a mixed population (Najafi et al., 2020) offers significant advantages for classification. However, we are working with a coarse-grained model of metapopulation comprising excitatory and inhibitory subpopulations. As a result, discrepancies with a real brain may occur (Roach et al., 2023; Kuan et al., 2024). In fact, besides the discrepancy in the ratio between inhibitory and excitatory subpopulations at a single vertex, the topological structure of the learned graph also does not follow assortative modularity, where nodes connect densely within their own community and sparsely to nodes outside their community. Instead, we obtain a fully connected weighted graph. In contrast to this topological structure, we observe that our adjecency matrix $\mathbf{A}$ remains asymmetric throughout all the analyses performed in this manuscript. Indeed, asymmetries in predicted communication efficiency reflect neurobiological concepts of functional hierarchy and correlate with directionality in resting-state effective connectivity, as analyzed using spectral dynamic causal modeling (Seguin et al., 2019; Patankar et al., 2020). Moreover, a principle of brain organization is that reciprocal connections between cortical areas are functionally asymmetric (Frässle et al., 2021).
@@
Table 1: Results of $\gamma$ over five different training runs for the MNIST and Fashion MNIST datasets. The numbers in parentheses represent the standard deviation of the mean and they refer to the last digits. Number of epochs for each single realization was set to $525$, with batch-size equal to $200$ for MNIST, while was set to $350$, with batch-size equal to $200$, for FASHION-MNIST.
@@
Table 2: Results of accuracy of our model (first column) and accuracy of MLP (second column), over five different training runs for the MNIST and Fashion MNIST datasets. The numbers in parentheses represent the standard deviation of the mean and they refer to the last digits. Number of epochs for each single realization was set to $525$, with batch-size equal to $200$ for MNIST, while was set to $350$, with batch-size equal to $200$, for FASHION-MNIST.
This model also offers a simple invertible property. By employing a straightforward transformation, $\tau=T-t$ (Marino and Macris, 2023), which effectively reverses the dynamics of the system, we can integrate the dynamics backwards, starting from the terminal position of the forward dynamics, and rebuild the initial condition from which we started with. The modified evolutionary law for this backward integration becomes $\dot{\mathbf{z}}=-G(\mathbf{z})$, with the temporal derivative now taken concerning the variable $\tau$. A simple cartoon is presented in Fig. 8.
Our analysis was performed on simple grayscale image datasets of size $28\times 28$. Such an analysis is computationally fast. However, testing colored images, such as those in the CIFAR-10 and TF-FLOWERS datasets, which have sizes $32\times 32\times 3$ and $224\times 224\times 3$ respectively, can be time-consuming due to the increased dimensionality of these images. To avoid such bottlenecks, we test our model with CNNs, which are biologically inspired. Additionally, we investigate whether the well-known concept of transfer learning (Hospedales et al., 2022) can be applied to our model. These analyses are presented in the following subsection.
@@ 5.2.1 Wilson-Cowan model for metapopulation in combination with a Convolutional Neural Network
Computational models validate intuitions about how a system works by providing a way to test those intuitions directly. They offer a means to explore new hypotheses in an ideal experiment. CNNs are an example. They came from pioneering research in neuroscience, which gave insights into the nature of visual processing in mammals including humans. In such a pioneering research, Hubel and Wiesel (1959) discovered two major cell types in the primary visual cortex (V1) of cats (see Fig. 9, Panel b)). The first type, the simple cells, respond to bars of light or dark when placed at specific spatial locations. The second type, complex cells, have less strict response profiles. These complex cells are likely receiving input from several simple cells, all with the same preferred orientation but with slightly different preferred locations. From such observations, Fukushima developed the Neocognitron (Fukushima, 1980), a precursor to modern CNN. This computational model contains two main cell types: the S-cells and the C-cells. The S-cells are named after simple cells and replicate their basic features on a plane, as well as the C-cells (named after complex cells) that are nonlinear function of several S-cells coming from the same plane but at different locations. After a plane of simple and complex cells representing the basic computations of V1, the Neocognitron simply repeats the process again. From these ideas, finally CNNs arose. A CNN is a type of deep learning model (Li et al., 2022) specifically designed for processing structured grid data, like images. CNNs automatically and adaptively learn spatial hierarchies of features from input images through convolutional layers, pooling layers, and fully connected layers. For more on architectures and applications of CNNs, we suggest the reviews (Lindsay, 2021; Zhao et al., 2024).
@@
CNNs can mimic and simulate the representation of visual information along the ventral stream (see Fig. 9, Panel a)). In particular, the activity of the artificial units of CNNs predicts the activity of real neurons in animals, with very high accuracy. For example, in (Yamins et al., 2014) the authors showed the existence of a strong correlation between a biologically plausible hierarchical neural network model’s categorization performance and its ability to predict individual inferior temporal (IT) neural unit response data. Furthermore, the activity of units from the last layer of the network best predicted IT activity and the penultimate layer best predicted V4, in visual cortex. Many other works have been developed in this direction, demonstrating that several deep CNN architectures exhibit similar performance to human and monkey object classification (Lindsay, 2021; Rajalingham et al., 2018).
@@
Figure 9: Description of biological and artificial object recognition. The figure illustrates a schematic representation of how object recognition and classification, in the brain and artificially, should work. Panel a) simply describes the position of the visual cortex. The visual cortex (grey, purple and green) is the primary cortical region of the brain that receives, integrates, and processes visual information relayed from the retinas. It is in the occipital lobe of the primary cerebral cortex, which is in the most posterior region of the brain. The visual cortex divides into five different areas (V1 to V5) based on function and structure. In the figure, only the positions of V1, V2, and V4 are presented for the sake of simplicity. Inferior Temporal (IT) is the cerebral cortex on the inferior convexity of the temporal lobe in primates including humans. It is crucial for visual object recognition and is considered to be the final stage in the ventral cortical visual system (grey and purple). The ventral stream transforms visual inputs into perceptual representations that embody the enduring characteristics of objects and their spatial relations. The ventral stream begins with V1, goes through visual area V2, then through visual area V4, and to the inferior temporal cortex. The ventral stream, is associated with form recognition, object representation and storage of long-term memory (Schneider, 1969; Milner and Goodale, 2008). The dorsal stream (grey and green) begins with V1, goes through area V2, then to the dorsomedial area and middle temporal area and to the posterior parietal cortex. The dorsal stream’s job is to mediate the visual control of skilled actions, such as reaching and grasping, directed at objects in the world. To do this, the dorsal stream needs to register visual information about the goal object on a moment-to-moment basis, transforming this information into the appropriate coordinates for the effector being used (Schneider, 1969; Milner and Goodale, 2008). Panel b) shows the relationship between components of the visual system (Biological) and the base operations of a CNN (Artificial). Biological: cartoon of simple cells (blue) and complex cells (green) (Hubel and Wiesel, 1959). Simple cells have preferred location in the image (dashed ovals). Complex cells receive input from many simple cells and thus have more spatially invariant responses. Artificial: the first convolutional layer (blue) is produced by applying a small filter (square box) to every location in the image. Such operation creates a collection of feature maps. For each feature map, for example, the maximal activation in the square blue box can be taken. Such an operation downsamples the image and leads to a complex cell-like plane (green). Panel c) illustrates a sketch of the CNN model that we use. It has a series of convolutional and pooling layers. It terminates with a set of dense layers that are connected to a Wilson-Cowan model for metapopulation. Given an input, the CNN feeds forward through all its layers to arrive at our model, which iteratively converges to the planted stationary state.
@@
Such a new model could, in principle, retain all the validated aspects of a visual system provided by the CNN part and then, at the end of the ventral cortical visual system (i.e., the IT), incorporate a recurrent network—our Wilson-Cowan model for metapopulation—that contains some long-term memories, which we define as our planted attractors. The IT region, indeed, is a brain region where visual perception meets memory and imagery (Miyashita, 1993). Long-term memory refers to the brain’s process of taking information from short-term memory and creating long-lasting memories (She et al., 2024). These memories are stable and easily accessible, like our planted attractors. Moreover, long-term memories can include information related to activities learned through practice and repetition, similar to how a learning algorithm operates. However, validating such a computational model is extremely challenging. We must begin, therefore, with the initial steps of implementing the model and analyzing its performance on various visual classification tasks. We defer in-depth analyses for validating the model with neurobiological data to future publications.
@@
  • a
    (Byerly et al., 2021)
  • b
    (Tanveer et al., 2021)
  • c
    SOTA obtained by using transformer architecture (Dosovitskiy et al., 2021). The best VGG-16 (Simonyan and Zisserman, 2015) has an accuracy of 0.93(Giuste and Vizcarra, 2020).
  • d
    SOTA obtained by using transformer architecture(Kolesnikov et al., 2020).
@@
For both datasets, the $\gamma$ parameter appears to remain close to the accepted values for a biological model (see Tab. 4). Indeed, the populations of excitatory neurons in all cortical areas should be around $70-80\%$ (DeFelipe and Fariñas, 1992), so our optimal values fell within the right range for the ratio between inhibitory and excitatory neurons.

### v2

@@ 3 Enforcing Linear Stability
the maximum value between the two real part of (7), i.e., $\max\{\text{Re}(s^{(+)}),\text{Re}(s^{(-)})\}$, must be taken. If such a value is negative, then the associated fixed point is asymptotically stable [88].
To be explicit, the elements of $[\mathbf{J}]_{ii}(\vec{\mathbf{z}}^{*})$ have the following expressions:
@@ 4 Training
Obtained the stability criteria for our attractors, we can construct a statistical learning model for our classification task. In such a statistical learning model only the components of the matrix $\Phi$, which are not the embedded eigenvectors, the non-zero eigenvalues in $\Lambda$, and the $\gamma$ parameter can be learned. In other words, the matrix $\Phi$ should be regarded as a matrix where the first $K$ columns, identifying the embedded attractors of the system’s dynamics, are not to be learned, just as the first $K$ eigenvalues of the matrix $\Lambda$, which are fixed to be zero by construction. This approach ensures that the planted attractors guide the dynamics, letting the system to converge on selected attractors once the training has been completed. As an initial condition for the non-embedded eigenvectors, we opt for orthogonal random eigenvectors [5]. As an initial condition for $\gamma$ we opt for a value equal to $0.25$ (as described in Section 2), while for the trainable eigenvalues we opt to initialize them with a normal distribution with mean equal to $-\sqrt{\mathcal{N}}$ and variance equal to $1$. The condition on the trainable eigenvalues satisfies completely the initial stability condition (see Fig. 3). Without loss of generality, we fix the non trainable parameters of the model, as described in the caption of Fig. 3 and used in [99].
@@
Figure 4: Building the model. The figure illustrates the composition of the dataset and the method for introducing stable attractors into the dynamics. Panel a) displays the labels, images, and targets used in our model. The first row shows the labels of the MNIST image ($\mathcal{N}=784$) examples found in the second row. The third row presents the ten target images we created to serve as attractors for the dynamics. Panel b) shows the creation of an eigenvector of matrix $\mathbf{A}$. The image is composed solely of the excitatory stationary parts of the two fixed points in the Wilson-Cowan model. Each black pixel corresponds to a $\overline{x}^{(1)}$, while each white pixel corresponds to a $\overline{x}^{(2)}$. In general, to create $K$ target vectors for a dataset, we generate $K$ vectors of size $\mathcal{N}$, where $\frac{\mathcal{N}}{K+2}$ components are set to $\overline{x}^{(1)}$ and the remaining components to $\overline{x}^{(2)}$. Each unique target vector is constructed by setting the components from $k\frac{\mathcal{N}}{K+2}$ to $((k+1)\frac{\mathcal{N}}{K+2})-1$ to $\overline{x}^{(1)}$, with all other components set to $\overline{x}^{(2)}$, with $k=0,\dots,K-1$ that identifies the label of a single class. Panel c) illustrates the positions of the fixed eigenvectors and eigenvalues (in red) and the trainable ones (in green). Recall that the matrix $\mathbf{A}$ is diagonalizable by definition.
@@
The input datum $\mathbf{z}^{(j)}$ represents, for our dynamical system defined in (1), an initial condition of the dynamics, i.e., $\mathbf{z}^{(j)}(t=0)$. From such initial condition, we evolve the dynamical system for a sufficiently large time $T$. The evolution is performed using Euler’s algorithm, discretizing time into steps of $\Delta t=0.1$, unless explicitly specified otherwise. At each iteration, denoted with the index $n$, the system’s state $\mathbf{z}^{(j)}_{n}$ is updated according to the equation: $\mathbf{z}^{(j)}_{n+1}=\mathbf{z}^{(j)}_{n}+G(\mathbf{z}^{(j)}_{n})\Delta t$, where $G(\mathbf{z}^{(j)})$ is the system’s flux, defined in (1). The total number of iterations, $n_{max}$, is determined a priori based on the desired simulation time $T$ and the chosen time step $\Delta t$ through the relation $n_{max}=\lfloor\frac{T}{\Delta t}\rfloor.$ This approach allowed us to efficiently simulate the system’s long-term behavior by iteratively applying Euler’s update rule. Once the system reaches the final iteration, we use the value of the solution to optimize the loss function.
To minimize the loss function in (9), we specifically employ Accelerated Stochastic Gradient Descent using the Adam optimizer [46], with a learning rate of $0.1$, unless explicitly specified otherwise. The learning rate, the number of epochs needed to achieve optimal minimization of the loss function, the mini-batch size used in the Adam algorithm, and the time $T$ required to reach the stationary state have been optimized empirically (see A).
@@
Figure 5: Dynamics of the model. The figure illustrates a schematic representation of our Neural Mass Network Model. Panel a) displays an image from the MNIST ($\mathcal{N}=784$) dataset along with its associated target. The image features a black pixel with a red border, which we track throughout its dynamics. Panel b) shows this dynamic process. The equations are highlighted in red to emphasize that we are focusing on that single pixel. This pixel is depicted as a red sphere moving within a double well potential. Our model, we recall, is configured to exhibit a bistable profile. Initially, the vector $\vec{\zeta}_{i}$ is set to represent the black pixel for both subpopulations of neurons. As the dynamics evolve according to our model’s equations, the pixel explores the bistable profile. It does not stop at the closest minimum; instead, it reaches the correct minimum designated for classification, thanks to the learned coupling in the matrix $\mathbf{A}$. Panel c) illustrates the overall dynamics of the entire image. Starting from an initial condition—the image we want to classify—the system uses the equations in (1) to achieve the final state, which is our stable attractor. The precision cutoff for this analysis has been set to the seventh decimal digit.
@@
$\delta(\arg\max\big{(}\vec{q}^{(j)}\big{)})$ is, therefore, a binary random variable that allows us to estimate the probability of success of our algorithm.
The computational complexity of our model, after the training process is complete, is quantified as $O(\mathcal{N}^{2})$, resulting from simple matrix-vector multiplication during the integration of the dynamics, where $\mathcal{N}$ represents the size of the input.
@@ 5 Experimental Validations
    Model Fine-Tuning with Pre-Trained Weights: The second subsection explores the test set performance of a Wilson-Cowan model for metapopulation in combination with convolutional neural networks [54] or a Pre-Trained Transformer [92]. In the first case we will analyse the performance of the combination across MNIST, Fashion-MNIST, CIFAR10, and TF-FLOWERS. While in the second case, we perform a classification for the sentiment analysis, i.e. IMDB dataset [73]. For details on these datasets, see A.
@@ 5.1 Training from Scratch
A critical distinction between the top panel of Fig. 6 and Fig. 3 lies in the value of the parameter $\gamma$, which plays a key role in governing the time scales within the standard Wilson-Cowan model. While traditionally it is set to $0.25$, our work explores the impact of optimizing $\gamma$ during the training process. In Fig. 3, $\gamma$ begins with its typical value of $0.25$. As stated before, a population in each vertex is composed by $80\%$ of excitatory and $20\%$ inhibitory neurons. However, during training, it undergoes optimization and converges to an optimal value of $0.839\pm 0.031$ for the MNIST dataset ( for Fashion MNIST results see Tab. 1). In this case, we have that a population in a vertex is composed by $54\%$ of excitatory and $46\%$ inhibitory neurons. These values suggest that, for our particular Neural Mass Network Model, a mixed population offers significant advantages for classification. However, we are working with a coarse-grained model of metapopulation comprising excitatory and inhibitory subpopulations. As a result, discrepancies with a real brain may occur. In fact, besides the discrepancy in the ratio between inhibitory and excitatory subpopulations at a single vertex, the topological structure of the learned graph also does not follow assortative modularity, where nodes connect densely within their own community and sparsely to nodes outside their community. Instead, we obtain a fully connected weighted graph. In contrast to this topological structure, we observe that our adjecency matrix $\mathbf{A}$ remains asymmetric throughout all the analyses performed in this manuscript. Indeed, asymmetries in predicted communication efficiency reflect neurobiological concepts of functional hierarchy and correlate with directionality in resting-state effective connectivity, as analyzed using spectral dynamic causal modeling [74]. Moreover, a principle of brain organization is that reciprocal connections between cortical areas are functionally asymmetric [31].
@@
Table 1: Results of $\gamma$ over five different training runs for the MNIST and Fashion MNIST datasets ($\mathcal{N}=784$). The numbers in parentheses represent the standard deviation of the mean and they refer to the last digits. Number of epochs for each single realization was set to $525$, with batch-size equal to $200$ for MNIST, while was set to $350$, with batch-size equal to $200$, for FASHION-MNIST.
@@
Table 2: Results of accuracy of our model (first column) and accuracy of MLP (second column), over five different training runs for the MNIST and Fashion MNIST datasets ($\mathcal{N}=784$). The numbers in parentheses represent the standard deviation of the mean and they refer to the last digits. Number of epochs for each single realization was set to $525$, with batch-size equal to $200$ for MNIST, while was set to $350$, with batch-size equal to $200$, for FASHION-MNIST.
To compare our method with other approaches in the current literature, we reviewed studies on the MNIST and Fashion-MNIST datasets that utilized various techniques with attractors in their dynamics, as in Hopfield networks that are Recurrent Neural Networks and biological inspired. However, these methods, at best, achieved performance metrics below $70\%$ for the MNIST dataset[8, 53], and $63\%$ for the Fashion-MNIST dataset [30]. Even with a supervised approach that transforms Hebb’s rule into a genuine learning rule, the accuracy only reached $94\%$ for MNIST and $84\%$ for Fashion-MNIST [3].
Additionally, we compared our results with those obtained by Krotov and Hopfield in their work [52]. In this study, the authors demonstrated that the Hopfield network can be modified by using a non-linear (e.g., polynomial) function in the Hamiltonian, allowing it to store a polynomial number of memories, rather than just a linear one. These memories are stable attractors of the discrete dynamics. Furthermore, the authors showed that these modified networks, known as Dense Associative Memory (DAM) networks, can also function as classifiers when a continuous non-linear function is applied in the asynchronous dynamics and a stochastic gradient descent is used to optimize the weights.
For instance, DAMs achieve a test error of $1.6\%$ on the MNIST dataset, while our Wilson-Cowan model for metapopulation achieves a test error of $1.9\%$. In other words, our model performs on par with DAMs.
This model also offers a simple invertible property. By employing a straightforward transformation, $\tau=T-t$ [60], which effectively reverses the dynamics of the system, we can integrate the dynamics backwards, starting from the terminal position of the forward dynamics, and rebuild the initial condition from which we started with. The modified evolutionary law for this backward integration becomes $\dot{\mathbf{z}}=-G(\mathbf{z})$, with the temporal derivative now taken concerning the variable $\tau$. A simple cartoon is presented in Fig. 8.
Our analysis was performed on simple grayscale image datasets of size $28\times 28$. Such an analysis is computationally fast, as showed in D for different image sizes. We performed also ablation experiments on our model, and such experiments are presented in E. However, testing colored images, such as those in the CIFAR-10 and TF-FLOWERS datasets, which have sizes $32\times 32\times 3$ and $224\times 224\times 3$ respectively, can be time-consuming due to the increased dimensionality of these images. To avoid such bottlenecks, we test our model with CNNs, which are biologically inspired. Additionally, we investigate whether the well-known concept of transfer learning [38] can be applied to our model. These analyses are presented in the following subsection.
@@ 5.2.1 Wilson-Cowan model for metapopulation in combination with a Convolutional Neural Network
Computational models validate intuitions about how a system works by providing a way to test those intuitions directly. They offer a means to explore new hypotheses in an ideal experiment. CNNs are an example. They came from pioneering research in neuroscience, which gave insights into the nature of visual processing in mammals including humans. In such a pioneering research, [41] discovered two major cell types in the primary visual cortex (V1) of cats (see Fig. 9, Panel b)). The first type, the simple cells, respond to bars of light or dark when placed at specific spatial locations. The second type, complex cells, have less strict response profiles. These complex cells are likely receiving input from several simple cells, all with the same preferred orientation but with slightly different preferred locations. From such observations, Fukushima developed the Neocognitron [32], a precursor to modern CNN. This computational model contains two main cell types: the S-cells and the C-cells. The S-cells are named after simple cells and replicate their basic features on a plane, as well as the C-cells (named after complex cells) that are nonlinear function of several S-cells coming from the same plane but at different locations. After a plane of simple and complex cells representing the basic computations of V1, the Neocognitron simply repeats the process again. From these ideas, finally CNNs arose. A CNN is a type of deep learning model [54] specifically designed for processing structured grid data, like images. CNNs automatically and adaptively learn spatial hierarchies of features from input images through convolutional layers, pooling layers, and fully connected layers. For more on architectures and applications of CNNs, we suggest the reviews [55, 100].
@@
CNNs can mimic and simulate the representation of visual information along the ventral stream (see Fig. 9, Panel a)). In particular, the activity of the artificial units of CNNs predicts the activity of real neurons in animals, with very high accuracy. For example, in [98] the authors showed the existence of a strong correlation between a biologically plausible hierarchical neural network model’s categorization performance and its ability to predict individual inferior temporal (IT) neural unit response data. Furthermore, the activity of units from the last layer of the network best predicted IT activity and the penultimate layer best predicted V4, in visual cortex. Many other works have been developed in this direction, demonstrating that several deep CNN architectures exhibit similar performance to human and monkey object classification [55, 79].
@@
Figure 9: Description of biological and artificial object recognition. The figure illustrates a schematic representation of how object recognition and classification, in the brain and artificially, should work. Panel a) simply describes the position of the visual cortex (see C). Panel b) shows the relationship between components of the visual system (Biological) and the base operations of a CNN (Artificial). Biological: cartoon of simple cells (blue) and complex cells (green) [41]. Simple cells have preferred location in the image (dashed ovals). Complex cells receive input from many simple cells and thus have more spatially invariant responses. Artificial: the first convolutional layer (blue) is produced by applying a small filter (square box) to every location in the image. Such operation creates a collection of feature maps. For each feature map, for example, the maximal activation in the square blue box can be taken. Such an operation downsamples the image and leads to a complex cell-like plane (green). Panel c) illustrates a sketch of the CNN model that we use. It has a series of convolutional and pooling layers. It terminates with a set of dense layers that are connected to a Wilson-Cowan model for metapopulation ($\mathcal{N}=784$). Given an input, the CNN feeds forward through all its layers to arrive at our model, which iteratively converges to the planted stationary state.
@@
Such a new model could, in principle, retain all the validated aspects of a visual system provided by the CNN part and then, at the end of the ventral cortical visual system (i.e., the IT), incorporate a recurrent network—our Wilson-Cowan model for metapopulation—that contains some long-term memories, which we define as our planted attractors. The IT region, indeed, is a brain region where visual perception meets memory and imagery [68]. Long-term memory refers to the brain’s process of taking information from short-term memory and creating long-lasting memories [85]. These memories are stable and easily accessible, like our planted attractors. Moreover, long-term memories can include information related to activities learned through practice and repetition, similar to how a learning algorithm operates. However, validating such a computational model is extremely challenging. We must begin, therefore, with the initial steps of implementing the model and analyzing its performance on various visual classification tasks. We defer in-depth analyses for validating the model with neurobiological data to future publications.
@@
For both datasets, the $\gamma$ parameter appears to remain close to the accepted values for a biological model (see Tab. 4). Indeed, the populations of excitatory neurons in all cortical areas should be around $70-80\%$ [21], so our optimal values fell within the right range for the ratio between inhibitory and excitatory neurons.